Publications: Robert L. Marcialis

Bell, B.M., Hall, D.S., and Marcialis, R.L. (1983) The photometric period of SZ Pictoris. Information Bulletin on Variable Stars No. 2272.

Bolewski, C., Lebofsky, N., Lebofsky, L., and Marcialis, R. (1992) Our solar system: Booklet, Grades K–3. (Our Solar System Series, Arizona Board of Regents), 27 pp.

Bolewski, C., Lebofsky, N., Lebofsky, L., and Marcialis, R. (1992) Our solar system: Booklet, Grades 2–5. (Our Solar System Series, Arizona Board of Regents), 40 pp.

Bolewski, C., Lebofsky, N., Lebofsky, L., and Marcialis, R. (1992) Our solar system: teacher’s guide and activities. (Our Solar System Series, Arizona Board of Regents), 46 pp.

Bolewski, C., Lebofsky, N., Lebofsky, L., and Marcialis, R. (1992) Nuestro sistema solar: Booklet, Grades K–3. (Our Solar System Series, Arizona Board of Regents), 27 pp.

Bolewski, C., Lebofsky, N., Lebofsky, L., and Marcialis, R. (1992) Nuestro sistema solar: Booklet, Grades 2–5. (Our Solar System Series, Arizona Board of Regents), 40 pp.

Boss, B.J., Smith, P.H., Tanner, R., Reynolds, R., and Marcialis, R. (2008) 2007 Phoenix Mars Scout Mission and Mars Surveyor 2001 Robotic Arm Camera (RAC) Calibration Report. Archived in the NASA Planetary Data System (PDS), http://pds.nasa.gov, 146 pp.

Boynton, W.V., Feldman, W.C., Squyres, S.W., Prettyman, T., Brückner, J., Evans, L.G., Reedy, R.C., Starr, R., Arnold, J.R., Drake, D.M., Englert, P.A.J., Metzger, A.E., Mitrofanov, I., Trombka, J.I., d’Uston, C., Wänke, H., Gasnault, O., Hamara, D.K., Janes, D.M., Marcialis, R.L., Maurice, S., Mikheeva, I., Taylor, G.J., Tokar, R., and Shinohara, C. (2002) Distribution of hydrogen in the near-surface of Mars: evidence for sub-surface ice deposits. Science 297, 81–85.

Boynton, W.V., Feldman, W.C., Mitrofanov, I., and Team GRS. (2002) Near-surface ice on Mars: early results from the 2001 Mars Odyssey Gamma Ray Spectrometer (GRS) instrument suite. Eos 83, no. 19, S222 (Abstract).

Boynton, W.V., Hamara, D.K., Feldman, W.C., Prettyman, T.H., and Mars Odyssey Gamma Ray Spectrometer Team (2002) Subsurface ice content in the north polar region of Mars. Bull. Amer. Astron. Soc. 34, no. 3, 852 (Abstract).

Buie, M.W., Olkin, C., McDonald, S., Ford, C., Foust, J., Sopata, L., Elliot, J.L., Marcialis, R., Hill, R., Bus, S.J., and Meserole, R., Wasserman, L., Spencer, J., Dunham, E., Young, J. Jarrett, T.H., Beichman, C.A., Herter, T., Telesco, C., Schulz, B., Campins, H., and Osip, D. (1993) (2060) Chiron IAU Circular No. 5898.

Buie, M.W., Wasserman, L., Millis, R.L., White, N.M., Nye, R., Dunham, E.W., Bosh, A.S., Stone, R., Hubbard, W.B., Hill, R., Dunham, D., Fried, R., Klinglesmith, IV, D., Sanford, J., Schwaar, P., Maley, P., Owen, W., Benner, L., Rivkin, A.S., Spitale, J., Marcialis, R.L., and Lebofsky, L.A. (1997) Occultation of GSC 23450183 by (704) Interamnia. Bull. Amer. Astron. Soc. 29, 973 (Abstract).

Buie, M.W., Keeney, B.A., Strauss, R.H., Blank, T.E., Moore, J.G., Porter, S.B., Wasserman, L.H., Weryk, R.J., Levison, H.F., Olkin, C.B., Leiva, R., Bardecker, J.E., Brown, M.E., Brown, L.B., Collins, M.P., Davidson, H.M., Dunham, D.W., Dunham, J.B., Eaccarino, J.A., Finley, T.J., Fuller, L., Garcia, M.L., George, T., Getrost, K., T. Gialluca, M.T., Givot, R,M., Gupton, D., Hanna, W.H., Hergenrother, C.W., Hernandez, Y., Hill, B., Hinton, P.C., Holt, T.R., Howell, R.R., Jewell, J.L., Kamin, R.L., Kammer, J.A., Kareta, T., Kayl, G.J., Keller, J.M., Kenyon, D.A., Kester, II, S.R., Kidd, Jr., J.N., Lauer, T.R., Leung, C.W.S., Lorusso, Z.R., Lundgren, C.B., Magana, L.O., Maley, P.D., Marchis, F., Marcialis, R.L., McCandless, A.E., y McCrystal, D.J., McGraw, A.M., Miller, K.E., Mueller, B.E.A., Noonan, J.W., Olsen, A.M., Patton, A.R., Peluso, D.O., Person, M.J., Rigby, J.G., Rolfsmeier, A.D., Salmon, J.J., Samaniego, J., Sawyer, R.P., Schulz, D.M., Skrutskie, M.F., Smith, R.J.C., Spencer, J.R., Springmann, A., Stanbridge, D.R., Stoffel, T.J., Tamblyn, P., Tobias, B., Verbiscer, A.J., von Schalscha, M.P., Werts, H., and Zhang, Q. (2021) Size and shape of (11351) Leucus from five occultations. Planetary Sci. Jour. 2, no. 5, 202.

Buratti, B.J., Marcialis, R.L., and Dunbar, R.S. (1991) Photometry of the comet 2060 Chiron. Reports of the Planetary Geology and Geophysics Program—1991. NASA Technical Memorandum 4300, 408–409.

Buratti, B.J., Marcialis, R.L., Howell, E.S., and Nolan, M.C. (1991) CCD Photometry of 2060 Chiron, 1991 January. International Conference on Asteroids, Comets and Meteors, Flagstaff, AZ, 1991 June, p. 33 (Abstract).

Buratti, B.J., Marcialis, R.L., and Foust, J. (1992) An outburst on 2060 Chiron in January 1992. Eos 73, 324–325 (Abstract).

Buratti, B.J., Dunbar, R.S., Tedesco, E.F., Gibson, J., Marcialis, R., Wong, F., and Dobrovolskis, A. (1992) Pluto–Charon mutual eclipse events: CCD observations and modeling. Bull. Amer. Astron. Soc. 24, 962 (Abstract).

Buratti, B.J., Marcialis, R., Dobrovolskis, A., Dunbar, R.S., Tedesco, E., Gibson, J., Mosher, J., and Wong, F. (1993) Pluto–Charon mutual eclipses: CCD observations from Palomar Mountain. Pluto and Charon, Flagstaff, AZ, 1993 July 10–14 (Abstract).

Buratti, B.J., Dunbar, R.S., Tedesco, E.F., Gibson, J., Marcialis, R.L., Wong, F., and Bennett, S. (1995) Modeling Pluto–Charon mutual eclipse events II. CCD observations with the 60-inch telescope at Palomar Mountain. Astron. Jour. 110, 1405–1419.

Buratti, B.J., Kangas, M., Foust, J., and Marcialis, R. (1993) Observations of the comet-asteroid 2060 Chiron over a continuous 6-day period. Eos 74, 385 (Abstract).

Buratti, B., Marcialis, R., Foust, J., and Kangas, M. (1994) “The bursty nature of 2060 Chiron.” Paper given at Completing the inventory of the solar system, Flagstaff, AZ, 27–30 June 1994.

Bus, S.J., Buie, M.W., Schleicher, D.G., Hubbard, W.B., Marcialis, R.L., Hill, R., Wasserman, L.H., Spencer, J.W., Millis, R.L., Franz, O.G., Bosh, A.S., Dunham, E.W., Ford, C.H., Young, J.W., Elliot, J.L., Meserole, R., Olkin, C.B., McDonald, S.W., Foust, J.A., Sopata, L.M., and Bandyopadhyay, R.M. (1996) Stellar occultation by 2060 Chiron. Icarus 123, 478–490.

Chapman, C.R., Davis, D.R., Weidenschilling, S.J., Marcialis, R., Greenberg, R., Levy, D., and Vail, S. (1986) Photometry of large rapidly rotating main–belt asteroids. Bull. Amer. Astron. Soc. 18, 796 (Abstract).

Crowe, D.G., Smith, P.H., Weinberg, J.D., Chabot, N., Hoppa, G., Tanner, R., Herkenhoff, K., Shinohara, C., Britt, D.T., Burkland, M., Hviid, S., Gunnlaugsson, H.G., Goetz, W., Thomas, N., Marcialis, R., Friedman, T., Doose, L., Reynolds, R., and Wellman, J. (1996) Imager for Mars Pathfinder Calibration Report.

Drish, Jr., W.F., Harmon, R., Wild, W.J., and Marcialis, R.L. (1993) Pluto images generated via Matrix Lightcurve Inversion. Pluto and Charon, Flagstaff, AZ, 1993 July 10–14 (Abstract).

Drish, Jr., W.F., Harmon, R., and Marcialis, R.L., and Wild, W.J. (1995) Images of Pluto generated by matrix lightcurve inversion. Icarus 113, 360–386.

Drish, Jr., W.F., Wild, W.J., Harmon, R., and Marcialis, R.L. (1995) Applications of matrix lightcurve inversion: an image of Pluto. Proc. SPIE 2622, 27–35.

Elliot, J.L., Dunham, E.W., Olkin, C.B., Ford, C., Stone, R.P.S., McDonald, S.W., Bandyopadhyay, R.M., Gilmore, D.K., Rank, D.M., Temi, P., Lazzaro, D., Hubbard, W.B., Reitsema, H., Barucci, A., Barroso, J., Bus, S.J., Lopes, D.F., Buie, M., Sicardy, B., Marcialis, R., Foryta, D.W., and Kurtz, D. (1994) Occultation by (2060) Chiron. IAU Circular No. 5965.

Elliot, J.L., Olkin, C.B., Dunham, E.W., Ford, C., Gilmore, D.K., Rank, D.M., Temi, P., Kurtz, D., Bandyopadhyay, R.M., Bus, S.J., McDonald, S.W., Barroso, J., Lazzaro, D., Lopes, D.F., Barucci, A., Bosh, A.S., Buie, M.W., Millis, R.L., Wasserman L.H., Dahn, C.C., Stone, R.C., Foryta, D.W., Hubbard, W.B., Marcialis R.L., Reitsema, H., Sicardy B., and Stone, R.P.S. (1994) Material near the nucleus of 2060 Chiron from stellar occultation observations. Bull. Amer. Astron. Soc. 26, 1153 (Abstract).

Elliot, J., Olkin, C., Hammel, H., Foust, J., McDonald, S.W., Dunham, E., Stone, R., Mattei, M., Bosh, A., Dahn, C., Cooray, A., Young, L., Buie, M., Wasserman, L., Millis, R.L., Hill, R., Reitsema, H., Howell, R., Wild, W., Drish, B., Briggs, J., Rank, D., Gilmore, D., Holbrook, J., Hubbard, W., and Marcialis, R.L. (1995) Stellar occultations by Neptune I (Triton). IAU Circular No. 6215.

Elliot, J.L., Olkin, C.B., Dunham, E.W., Ford, C., Gilmore, D.K., Kurtz, D., Lazzaro, D., Rank, D.M., Temi, P., Bandyopadhyay, R.M., Barroso, J., Barucci, A., Bosh, A.S., Buie, M.W., Bus, S.J., Dahn, C.C., Foryta, D.W., Hubbard, W.B., Lopes, D.F., Marcialis, R.L., McDonald, S.W., Millis, R.L., Reitsema, H., Schleicher, D.G., Sicardy B., and Stone, R.P.S., and Wasserman, L.H. (1995) Jet-like features near the nucleus of Chiron. Nature 373, 46–49.

Fernández, Y.R., Lisse, C.M., Wellnitz, D.D., A’Hearn, M.F., Kundu, A., Buie, M.W., Dunham, E.W., Millis, R.L., Nye, R.A., Stansberry, J.A., Wasserman, L.H., Golden, M.E., Persson, M., Howell, R.R., Marcialis, R.L., Spitale, J.N., Hoffmann, W.F., Dayal, A., Deutsch, L.K., Faziom G.G., Hora, J.L., Hanner, M.S., Ressler, M., Kostiuk, T., and Livengood, T.A. (1997) Multiwavelength observations of the nucleus and coma of comet Hale–Bopp. Bull. Amer. Astron. Soc. 29, 1047 (Abstract).

Fernández, Y.R., Wellnitz, D.D., Buie, M.W., Dunham, E.W., Millis, R.L., Nye, R.A., Stansberry, J.A. Wasserman, L.H., A’Hearn, M.F., Lisse, C.M., Golden, M.E., Person, M.J., Howell, R.R., Marcialis, R.L., and Spitale, J.N. (1999) The inner coma and nucleus of comet Hale–Bopp: results from a stellar occultation. Icarus 140, 205–220.

Fink, U., Schultz, A., DiSanti, M., and Marcialis, R. (1985) CCD imaging and spectroscopy of Comet P/Giacobini–Zinner (1984e). Bull. Amer. Astron. Soc. 17, 686 (Abstract).

The Galilean Satellite Observers: Franklin, F., Africano, J., Allen, W., Aksnes, K., Birch, P., Blanco, C., Coulson, I., Goguen, J., Jones, R., Langhans, T., Marcialis, R., Millis, R., Nakamura, T., Rettig, D., Rowe, C., Sinton, W., Solheim, J.-E., Skillman, D., Walters, J., Wasserman, L., Weller, W., Westphall, J. (1991) An analysis of the 1985 observations of mutual phenomena of the Galilean satellites. Astron. Jour. 102, 806–815.

Gehrels, T., Marcialis, R.L., Means, D., Nelson, M.L., Rabinowitz, D., and Scotti, J.V. (1990) Observations of minor planets — 1990 KA. Minor Planet Circular 16, 528–529.

Greenberg, R., Croft, S.K., Eplee, G., Janes, D.M., Kargel, J., Lebofsky, L., Lunine, J., Marcialis, R., Melosh, H.J., Ojakangas, G., and Strom, R. (1988) Miranda. Uranus: A Conference and a Book, 1988 June, Pasadena, CA (Abstract).

Greenberg, R., Croft, S.K., Janes, D.M., Kargel, J.S., Lebofsky, L.A., Lunine, J.I., Marcialis, R.L., Melosh, H.J., Ojakangas, G.W., and Strom, R.G. (1991) “Miranda.” In Uranus, J.T. Bergstrahl, E.D. Miner, and M.S. Matthews, eds. (Tucson: Univ. Arizona Press), pp. 693–735.

Hardie, R.H., Marcialis, R.L., Wilson, J.W., and Furman, W.R. (1985) Astrometric observations of Pluto: 1965–81. Astron. Jour. 90, 2643.

Hilton, D., Eplee, E., Grinspoon, D., Hildebrand, A., Jones, T., Marcialis, R., Marley, M., Pope, S., Rizk, B., Schneider, N., Tyler, A., and Hunten, D.M. (1985) Construction of the ‘Mars Ball’ prototype exploration vehicle. Bull. Amer. Astron. Soc. 17, 697 (Abstract).

Howell, E., Marcialis, R., Cutri, R., Nolan, M., Lebofsky, L., and Sykes, M. (1992) 1992 AD. IAU Circular No. 5449.

Howell, E., Rivkin, A.S., Marcialis, R.L., Nolan, M.C., Grier, J.A., Bus, S.J., and Slivan, S.M. (1994) Spectral observations of near-earth asteroids. Bull. Amer. Astron. Soc. 26, 1172–1173 (Abstract).

Hubbard, W.B., Lebofsky, L.A., Hunten, D.M., Reitsema, H.J., Zellner, B.H., Goff, R., Marcialis, R., Sykes, M., Frecker, J., Sánchez, A., Ríos, H.M., and Izaguirre, M.M. (1985) Occultation diameter of 1 Ceres. Lunar & Planetary Sci. 16, 370–371.

Hubbard, W.B., Rieke, G.H., Rieke, M.J., Lebofsky, L.A., and Marcialis, R.L. (1987) A grazing occultation by Neptune on 4 May 1986.Icarus In preparation.

Hubbard, W.B., Rieke, G.H., Rieke, M.J., Marcialis, R., and Campins, H. (1987) Coincidence search for Neptunian ring arcs. Bull. Amer. Astron. Soc. 19, 885 (Abstract).

Hubbard, W.B., Rieke, G.H., Rieke, M.J., Hunten, D.M., Marcialis, R., Lebofsky, L.A., Nicholson, P.D., Campins, H., and Matthews, K. (1989) Coincidence searches for Neptunian ring arcs in 1987 and 1988. In preparation.

Hubbard, W.B., Porco, C., Hunten, D., Rieke, G., Rieke, M., Asphaug, E., Clark, R., Haemmerle, V., Haller, J., Holberg, J., Lebofsky, L., Marcialis, R., McCarthy, D., McLeod, B., Buie, M., Elias, J., Jewitt, D., Persson, E., Boroson, T., West, S., Landau, R., and Schuster, W. (1989) Preliminary results from the occultation of 28 Sgr by the Saturn system: Saturn. Bull. Amer. Astron. Soc. 21, 951 (Abstract).

Hubbard, W.B., Porco, C., Clark, R., Turtle, E., Haemmerle, V., Hunten, D., Rieke, G., Rieke, M., Haller, J., Holberg, J., Lebofsky, L., Marcialis, R., McCarthy, D., McLeod, B., Buie, M., Elias, J., Jewitt, D., Persson, E., Boroson, T., West, S., Landau, R., and Carrasco, L. (1990) Saturn pole position and ring radius scale from 28 Sgr occultation. Bull. Amer. Astron. Soc. 22, 1041 (Abstract).

Hubbard, W.B., Porco, C.C., Hunten, D.M., Rieke, G.H., Rieke, M.J., McCarthy, D.W., Haemmerle, V., Clark, R., Turtle, E.P., Haller, J., McLeod, B., Lebofsky, L.A., Marcialis, R., Holberg, J., Landau, R., Carrasco, L., Elias, J., Buie, M.W., Jewitt, D., Persson, S.E., Boroson, T., West, S., and Mink, D.J. (1993) The occultation of 28 Sgr by Saturn: Saturn pole position and astrometry. Icarus 103, 215–234.

Hubbard, W.B., Porco, C.C., Hunten, D.M., Rieke, G.H., Rieke, M.J., McCarthy, D.W., Haemmerle, V., Haller, J., McLeod, B., Lebofsky, L.A., Marcialis, R., Holberg, J., Landau, R., Carrasco, L., Elias, J., Buie, M.W., Persson, S.E., Boroson, T., West, S., and French, R.G. (1993) Saturn’s mesospheric temperature from 28 Sgr occultations. Bull. Amer. Astron. Soc. 26, 1109 (Abstract).

Hubbard, W.B., Porco, C.C., Hunten, D.M., Rieke, G.H., Rieke, M.J., McCarthy, D.W., Haemmerle, V., Haller, J., McLeod, B., Lebofsky, L.A., Marcialis, R., Holberg, J., Landau, R., Carrasco, L., Elias, J., Buie, M.W., Dunham, E.W., Persson, S.E., Boroson, T., West, S., French, R.G., Harrington, J., Elliot, J.L., Forrest, W.J., Pipher, J.L., Stover, R.J., Brahic, A., and Grenier, I. (1997) Structure of Saturn’s mesosphere from the 28 Sgr occultations. Icarus 130, 404–425.

Keller, H.U., Hartwig, H., Kramm, R., Koschny, D., Markiewicz, W.J., Thomas, N., Fernandes, M., Smith, P.H., Reynolds, R., Lemmon, M.T., Weinberg, J., Marcialis, R., Tanner, R., Bos, B.J., Oquest, C., and Paige, D. (2001) The MVACS Robotic Arm Camera. Jour. Geophys. Res. Planets 106, 17609–17621.

Larson, S. and Marcialis, R. (1992) (2060) Chiron. IAU Circular No. 5669.

Larson, S., Scotti, J., Nolan, M., Howell, E., Marcialis, R., A’Hearn, M., Wellnitz, D., Bessell, M., Brosch, N., Campins, H., Gilmore, A., Kervin, P., Africano, J., Levine, S., Schuster, W., Martin, P., Moreno, F., Molina, A., Kidger, M., Rabinowitz, D., Butner, H., Tapia, S., Vilas, F., and Doherty, P. (1994) The Comet Impact Network Experiment (CINE): Preliminary results. Eos 75, 403–404 (Abstract).

Larson, S., Scotti, J., Howell, E., Marcialis, R., Nolan, A’Hearn, M., Wellnitz, D., Campins, H., Rabinowitz, D., Butner, H., Vilas, F., Doherty, P., Bessell, M., Brosch, N., Liebowitz, E., Gilmore, A., Kervin, P., Africano, J., Levine, S., Schuster, W., Martin, P., Verveer, A., Moreno, F., Molina, A., and Kidger, M. (1994) The Comet Impact Network Experiment (CINE) Bull. Amer. Astron. Soc. 26, 1570 (Abstract).

Larson, S., Scotti, J., Howell, E., Marcialis, R., Nolan, M., Wiśniewski, W., A’Hearn, M., Wellnitz, D., Bessell, M., Brosch, N., Campins, H., Gilmore, A., Kervin, P., Africano, J., Levine, S., Schuster, W., Martin, P., Moreno, F., Molina, A., Kidger, M., Rabinowitz, D., Butner, H., Tapia, S., Vilas, F., and Doherty, P. (1994) “The Comet Impact Network Experiment (CINE): early results.” Paper given at Hypervelocity Impact Symposium, Santa Fe, NM, 1994 Oct 17.

Lemmon, M.T., Yingst, R.A., Reid, R.J., Smith, P.H., and the MVACS Team (1999) “SSI/RAC: Imagers for Mars Polar Lander.” Paper given at Fifth International Conference on Mars, Pasadena, CA, 1999 July 19-24.

Marcialis, R.L. (1983) A two–spot model for the surface of Pluto. Master’s Thesis, Vanderbilt University. Marcialis, R.L. (1984) A two–spot model for the surface of Pluto. Bull. Amer. Astron. Soc. 16, 651

(Abstract). Marcialis, R.L. (1984) Single–Line Spectroscopic Binary. Contribution #41–03041 to the Hewlett-

Packard User’s Library. Marcialis, R.L. (1985) Spectrophotometry of Pluto–Charon mutual events. Paper given at the conference

Pluto at Perihelion, JPL, 1985 September 25 (abstract). Marcialis, R.L. (1985) Topographic relaxation on ice–covered worlds: application to Pluto. Bull. Amer.

Astron. Soc. 17, 715 (Abstract). Marcialis, R.L. (1985) Observations of mutual phenomena of the Galilean satellites in 1985–1986. Bull.

Amer. Astron. Soc. 18, 639 (Abstract). Marcialis, R.L. (1986) Photometry of Pluto–Charon mutual events in 1986. Bull. Amer. Astron. Soc. 18,

821–822 (Abstract). Marcialis, R. and Greenberg, R. (1987) Warming of Miranda during chaotic rotation. Nature 328,

227–228. Marcialis, R. and Greenberg, R. (1987) Warming of Miranda during chaotic rotation. Bull. Amer.

Astron. Soc. 19, 821 (Abstract). Marcialis, R.L., Rieke, G.R., and Lebofsky, L.A. (1987) Surface composition of Charon: tentative

identification of water ice. Symposium on the Origin and Evolution of Planetary and Satellite Atmospheres, Tucson, AZ, 1987 March 10–14 (Abstract).

Marcialis, R.L., Rieke, G.R., and Lebofsky, L.A. (1987) The surface composition of Charon: tentative

identification of water ice. Science 237, 1349–1351. Marcialis, R.L. (1987) The surface composition of Charon: tentative identification of water ice. Bull.

Amer. Astron. Soc. 19, 859 (Abstract). Marcialis, R.L. (1988) A two–spot albedo model for the surface of Pluto. Astron. Jour. 95, 941–947. Marcialis, R. and Fink, U. (1987) Spectrophotometry of Oberon. Bull. Amer. Astron. Soc. 19, 819

(Abstract). Marcialis, R.L. (1988) Spotting Pluto. Sky and Tel. 75, 6 (Letter to Editor). Marcialis, R.L. and Lebofsky, L.A. (1988) CVF spectrophotometry of Pluto and Triton. Bull. Amer.

Astron. Soc. 20, 807 (Abstract). Marcialis, R.L. (1988) The albedos of Pluto and Charon: wavelength dependence. Bull. Amer. Astron.

Soc. 20, 807 (Abstract). Marcialis, R.L. (1988) The discovery of Charon: happy accident or timely find? Bull. Amer. Astron. Soc.

20, 840 (Abstract). Marcialis, R.L. (1989) The discovery of Charon: happy accident or timely find? Jour. Brit. Astron. Assoc.

99, 27–29. Marcialis, R.L. and Lebofsky, L.A. (1989) Pitfalls to direct comparison of cometary and asteroidal

light curves. Submitted to Astrophys. Jour.

Marcialis, R.L. (1989) Topographic relaxation on ice–covered worlds: application to Pluto II. Eos 70, 386 (Abstract).

Marcialis, R.L. (1989) Topographic relaxation on ice–covered worlds: application to Pluto II. Submitted to Geophys. Res. Letters

Marcialis, R.L. (1989) CCD observations of 2060 Chiron, 1984–1985. Bull. Amer. Astron. Soc. 21, 965 (Abstract).

Marcialis, R.L. (1989) The New Pluto. Invited contribution to a NASA Planetary Astronomy Program brochure summarizing recent ground-based results regarding the Pluto–Charon system. (Submitted 13 February 1989).

Marcialis, R.L. (1990) The Pluto–Charon system as revealed during the mutual events. Ph.D. Thesis, University of Arizona.

Marcialis, R.L. (1990) The Pluto–Charon system as revealed during the mutual events. Pub. Astron. Soc. Pacific 102, 1224 (Abstract).

Marcialis, R.L. (1990) Compositional diversity in the Pluto–Charon system. Bull. Amer. Astron. Soc. 22, 1128 (Abstract).

Marcialis, R.L. and Lebofsky, L.A. (1991) CVF spectrophotometry of Pluto: correlation of composition with albedo. Icarus 89, 255–263.

Marcialis, R.L. and Buratti, B.J. (1991) CCD photometry of 2060 Chiron in 1985 and 1991. Bull. Amer. Astron. Soc. 23, 1155 (Abstract).

Marcialis, R.L. and Bus, S.J. (1991) The first Workshop on 2060 Chiron: What manner of beast is the Centaur? Bull. Amer. Astron. Soc. 23, 1157 (Abstract).

Marcialis, R.L., Lebofsky, L.A., DiSanti, M.A., Fink, U., Tedesco, E.F., and Africano, J. (1992) The albedos of Pluto and Charon: wavelength dependence. Astron. Jour. 103, 1389–1394.

Marcialis, R.L. and Buratti, B.J. (1993) CCD photometry of 2060 Chiron in 1985 and 1991. Icarus 104, 234–243.

Marcialis, R.L. (1993) The first half century of Pluto–Charon research. Pluto and Charon, Flagstaff, AZ, 1993 July 10–14 (Abstract).

Marcialis, R.L. (1993) Introducing PLUBIB: a Pluto–Charon bibliography database. Pluto and Charon, Flagstaff, AZ, 1993 July 10–14 (Abstract).

Marcialis, R.L. (1993) Correlation of composition with albedo on Pluto: past results and new problems. Pluto and Charon, Flagstaff, AZ, 1993 July 10–14 (Abstract).

Marcialis, R.L. and Larson, S.M. (1993) Deep CCD imaging and photometry of 2060 Chiron during the 1992–1993 apparition. Bull. Amer. Astron. Soc. 25, 1057 (Abstract).

Marcialis, R.L. (1993) Introducing PLUBIB: a Pluto–Charon bibliography database. Bull. Amer. Astron. Soc. 25, 1132 (Abstract).

Marcialis, R.L. (1993) Correlation of composition with albedo on Pluto: past results and new problems. Bull. Amer. Astron. Soc. 25, 1132 (Abstract).

Marcialis, R.L., Hubbard, W.B, Hill, R., Bus, S.J., Elliot, J.L., Olkin, C., McDonald, S., Foust, J., Sopata, L., Bandyopadhyay, R., Meserole, R., Buie, M.W., Spencer, J.R., Wasserman, L.H., Millis R., Dunham, E., Beichman, C.A., Jarrett, T.H., Young, J., Ford, C., and Herter, T. (1994) The 1993 Nov 07 occultation of Ch02 by 2060 Chiron. Bull. Amer. Astron. Soc. 26, 1153 (Abstract).

Marcialis, R. and Hill, R. (1994) Chiron credit. Sky and Tel. 87, 8 (Letter to Editor).

Marcialis, R.L. (1996) “The Pluto–Charon system.” In MacMillan Encyclopedia of the Earth Sciences, ed. E.J. Dasch (New York: Simon & Schuster-MacMillan Publishing), pp. 870–873.

Marcialis, R.L. (1997) “Introducing PLUBIB: a Pluto–Charon bibliography database.” In Pluto and Charon, S.A. Stern and D.J. Tholen, eds. (Tucson: Univ. Arizona Press), pp. 665–671..

(Current version on-line at: http://www.lpl.arizona.edu/ umpire/science/plubib home.html.)

Marcialis, R.L. (1997) “The first 50 years of Pluto–Charon research.” In Pluto and Charon, S.A. Stern and D.J. Tholen, eds. (Tucson: Univ. Arizona Press), pp. 27–83.

Marcialis, R.L. (1996) Pluto update: surface already mapped, photos prove to be model image. Physics Today 49, no. 7, 92–93 (Letter to editor).

Marcialis, R.L. (1996) A re-investigation of the 1965 April 29 stellar appulse by the Pluto–Charon system. Bull. Amer. Astron. Soc. 28, 1080 (Abstract).

Marcialis, R.L. (1998) A re-investigation of the 1965 April 29 stellar appulse by the Pluto–Charon system. Pub. Astron. Soc. Pacific. In preparation.

Marcialis, R.L. and Merline, W.J. (1998) The diameter of Pluto: a re-analysis of Kuiper’s disk meter measurements. Bull. Amer. Astron. Soc. 30, 1108 (Abstract).

Marcialis, R.L. (2002) “Pluto.” In MacMillan Encyclopedia of Space Sciences, Vol. 2 (Planetary Science and Astronomy), ed. P. Dasch (New York, NY: Macmillan Reference USA), pp. 147–151.

Marcialis, R.L. (2003) “The solar system.” In Encyclopedia of Optical Engineering, ed. R. Driggers (New York, NY: Marcel-Dekker Publishing), pp. Volume 3, 2591–2606.

Marcialis, R.L. (2003) “The solar system.” In Encyclopedia of Optical Engineering and Photonics, Second Edition, ed. C. Hoffman and R. Driggers (New York, NY: Taylor and Francis Publishing), pp. Volume V, 2591–2606.

Marcialis, R.L., Barlow, N.G., and Lebofsky, L.A. (2018) Nuclear nonsolution. Aerospace America 56, no. 4, 44–47.

Marcialis, R.L (2024) Signs of life. In “Aero Puzzler”. Aerospace America 62, no. 6, 8.

Marcialis, B., Vilas, F., Prato, L., and Hayden. L. (2020) Nadine G. Barlow (1958–2020). Bull. Amer. Astron. Soc. 52, no. 2, https://doi.org/10.3847/25c2cfeb.d309cee8.

Mechler, G., Croft, S.K., Hutson, M., and Marcialis, R. (1995) Planets and their Moons (National Audubon Society Pocket Guide, Chanticleer Press), 192 pp.

Mechler, G., Hutson, M., and Marcialis, R. (1995) The Sun and the Moon (National Audubon Society Pocket Guide, Chanticleer Press), 192 pp.

Millis, R., Wasserman, L., Franz, O., Hubbard, W., Lebofsky, L., Goff, R., Marcialis, R., Sykes, M., Frecker, J., Hunten, D., Reitsema, H., Zellner, B., Schneider, G., Ríos, M., Dunham, E., Klavetter, J., Meech, K., Oswalt, T., Rafert, J., A’Hearn, M., Osborn, W., Parker, D., Klemola, A., and Pirronen, J. (1985) The occultation diameter of Ceres. Bull. Amer. Astron. Soc. 17, 729 (Abstract).

Millis, R., Wasserman, L., Franz, O., Hubbard, W., Lebofsky, L., Goff, R., Marcialis, R., Sykes, M., Zellner, B., Reitsema, H., Schneider, G., Dunham, E., Klavetter, J., Meech, K., Oswalt, T., Rafert, J., Strother, E., Smith, J., Povenmire, H., Kornbluh, D., Reed, L., Izon, K., A’Hearn, M.F., Osborn, W., Parker, D., Douglas, W.T., Beish, J.D., Klemola, A., Ríos, M., Sánchez, A., Piironen, J., Mooney, M., Ireland, R.S., and Laibow, D. (1987) The size, shape, density, and albedo of Ceres from its occultation of BD+8^{∘}471. Icarus 72, 507–518.

Millis, R.L., Wasserman, L.H., Franz, O.G., Bowell, E., Nye, R., Thompson, D.T., White, N.M., Hubbard, W.B., Eplee Jr., R.E., Lebofsky, L.A., Marcialis, R.L., Greenberg, R.J., Hunten, D.M., Reitsema, H.J., Dunham, D.W., Maley, P., Klemola, A.R., Yeomans, D.K., and Bochen, Q. (1988) Observations of the 1987 occultation by 324 Bamberga. Asteroids II Conference, Tucson, AZ, 1988 March 8–11 (Abstract).

Millis, R.L., Wasserman, L.H., Franz, O.G., Bowell, E., Nye, R., Thompson, D.T., White, N.M., Hubbard, W.B., Eplee Jr., R.E., Lebofsky, L.A., Marcialis, R.L., Greenberg, R.J., Hunten, D.M., Reitsema, H.J., Bochen, Q., Dunham, D.W., Maley, P., Klemola, A.R., and Yeomans, D.K. (1988) Observations of the 8 December 1987 occultation of AG+40^{∘}0783 by 324 Bamberga. Astron. Jour. 98, 1094–1099.

Mitrofanov, I.G., Anfimov, D.S., Kozyrev, A.S., Litvak, M.L., Sanin, A.B., Tret’yakov, V.I., Boynton, W.V., Shinohara, C., Hamara, D., Saunders, R.S., H.E.N.D. Team, G.R.S. Team, and Odyssey Team. (2002) The stable and variable features of subsurface water on Mars: HEND data from Odyssey. Bull. Amer. Astron. Soc. 34, no. 3, 852 (Abstract).

Olkin, C.B., Elliot, J.L., Hammel, H.B., Cooray, A.R., McDonald, S.W., Foust, J.A., Bosh, A.S., Buie, M.W., Millis, R.L., Wasserman, L.H., Dunham, E.W., Young, L.A., Howell, R.R., Hubbard, W.B., Hill, R., Marcialis, R.L., McDonald, J.S., Rank, D.M., Holbrook, J.C., and Reitsema, H.J. (1997) The thermal structure of Triton’s atmosphere: results from the 1993 and 1995 occultations. Icarus 129, 178–201.

Porco, C., Hubbard, W., Hunten, D., Rieke, G., Rieke, M., Asphaug, E., Clark, R., Haemmerle, V., Haller, J., Holberg, J., Lebofsky, L., Marcialis, R., McCarthy, D., McLeod, B., Buie, M., Elias, J., Jewitt, D., Persson, E., Boroson, S., and West, S. (1989) Preliminary results from the occultation of 28 Sgr by the Saturn system: rings. Bull. Amer. Astron. Soc. 21, 928 (Abstract).

Reid, R.J., Smith, P.H., Lemmon, M., Tanner, R., Burkland, M., Wegryn, E., Weinberg, J., Marcialis, R., Britt, D.T., Thomas, N., Kramm, R., Dummel, A., Crowe, D.G., Bos, B.J., Bell III, J.F., Rueffer, P., Gliem, F., Johnson, J.R., Maki, J.N., Herkenhoff, K.E., and Singer, R.B. (1999) IMP image calibration. Jour. Geophys. Res. 104, no. E4, 8907–8926.

Reitsema, H.J., Hubbard, W.B., Hill, R., Marcialis, R., and Howell, R. (1995) WIRO observation of stellar occultation by Triton on 14 August 1995. Bull. Amer. Astron. Soc. 27, 1101 (Abstract).

Schultz, A., Fink, U., DiSanti, M., Buie, M., and Marcialis, R. (1985) CCD imaging and photometry of the development of P/Halley October 1984 – March 1985. Pub. Astron. Soc. Pacific 97, 903 (Abstract).

Scotti, J., Larson, S., Howell, E., Marcialis, R., Nolan, M., Wiśniewski, W. A’Hearn, M., Wellnitz, D., Campins, H., Rabinowitz, D., Butner, H., Tapia, S., Vilas, F., and Doherty, P., Bessell, M., Brosch, N., Liebowitz, E., Gilmore, A., Kervin, P., Africano, J., Levine, S., Schuster, W., Martin, R., Verveer, A., Moreno, F., Molina, A., Rutten, R., and Kidger, M. (1994) Preliminary results from CINE. Bull. Amer. Astron. Soc. 26, 1598 (Abstract).

Smith, P.H., Britt, D., Burkland, M., Chabot, N., Crowe, D.G., Friedman, T., Head, J., Hoppa, G., Marcialis, R.L., Reid, R., Reynolds, R., Shinohara, C., Singer, R., Tanner, R., Tomasko, M.G., Weinberg, J., Keller, H.U., Koschny, D., Kramm, R., Thomas, N., Gliem, F., Rueffer, P., Greeley, R., Sullivan, R., Knudsen, J.M., Madsen, B.M., Hviid, S., Gunnlaugsson, P., Soderblom, L., Gaddis, L., Ward, W., and Herkenhoff, K. (1996) The IMAGER for MARS PATHFINDER (IMP) experiment. Bull. Amer. Astron. Soc. 28, 1062–1063 (Abstract).

Smith, P.H., Reynolds, R., Weinberg, J., Friedman, T., Lemmon, M.T., Tanner, R., Reid, R.J., Marcialis., R.L., Bos, B.J., Oquest, C., Keller, H.U., Markiewicz, W.J., Kramm, R., Gliem, F., and Rueffer, P. (2001) The MVACS Surface Stereo Imager on Mars Polar Lander. Jour. Geophys. Res. Planets 106, 17589-17607.

Spanish Project Translation Board, Astronomical Society of the Pacific, (2002) El universo a sus pies; (A. Fraknoi and D. Schatz, eds.), Project ASTRO Translation Project, 490 pp.

Turtle, E., Porco, C., Hubbard, W., Haemmerle, V., Clark, R., Hunten, D., Rieke, G., Rieke, M., Haller, J., Holberg, J., Lebofsky, L., Marcialis, R., McCarthy, D., McLeod, B., Buie, M., Elias, J., Jewitt, D., Persson, E., Boroson, T., West, S., Landau, R., and Carrasco, L. (1990) The kinematics of Saturn’s major narrow rings from combined Voyager and ground-based data. Bull. Amer. Astron. Soc. 22, 1041 (Abstract).

Yingst, R.A., Smith, P.H., Lemmon, M.T., Marcialis, R.L., Rice, J.W., and Weinberg, J.D. (2001) DIRTCam in the desert: the Silver Lake field test of the Robotic Arm Camera. Jour. Geophys. Res. Planets 106, 7721–7732.

Total number of citations = 128.

Last edited: June 16, 2025